random231 (random231):

A binary operation defined as a*b=|a-b| is the binary operation associative?

random231 (random231):

@eliassaab can ya help me?

OpenStudy (anonymous):

why don't you check? :) is a*(b*c) always equal to (a*b)*c ?

random231 (random231):

uhh u hav ta prove it in exam okay? u cant tell the teacher why dontucheck it maam @PeterPan

OpenStudy (anonymous):

2*(-3*4)= 5 (2*(-3))*4=1 What can you conclude?

random231 (random231):

yeah i get it

random231 (random231):

so it has to be shown with an example?

random231 (random231):

okay thanks

OpenStudy (anonymous):

yw